OpenStudy (naythyngo):

will medal!!!! What is the probability of drawing three queens from a standard deck of cards, given that the first card drawn was a queen? Assume that the cards are not replaced. I'm not sure how to set up this problem

OpenStudy (kashmoneyjr):

Is your question asking what is the probability of drawing three queens on your next three turns?

OpenStudy (naythyngo):

yes its supposed to be a conditional probability question

OpenStudy (kashmoneyjr):

Ok so there are 52 cards in a deck, of those 4 queens

OpenStudy (kashmoneyjr):

youve already picked one queen, so there are now 51 cards and 3 queens

OpenStudy (kashmoneyjr):

what is the chance that you will pick a queen on your next turn?

OpenStudy (naythyngo):

I tried to set it up like 4/52 times 3/51times 3/50 but i didn't get an answer that matched the answer choices

OpenStudy (naythyngo):

2/50 not 3/50

OpenStudy (kashmoneyjr):

youre actually really close

OpenStudy (kashmoneyjr):

since you know that you picked a queen on your first pick, that is a given and has noting to do with your future probabilities except decrease the number of queens and number of cards in the deck by 1.

OpenStudy (kashmoneyjr):

so you can throw 4/52 out

OpenStudy (kashmoneyjr):

now you have to pick the three next cards: so, your probabilities will be, following the progression that you made: 3/51, 2/50, then 1/49

OpenStudy (kashmoneyjr):

and you multiply them together to get your answer. Does that make sense?

OpenStudy (naythyngo):

i tried it that way but I still didn't get a correct answer

OpenStudy (kashmoneyjr):

Hmm.. do you have the correct answer with you, by any chance?

OpenStudy (naythyngo):

no just the answer choices I can choose from but I think the answer is 1/425

OpenStudy (kashmoneyjr):

Actually, looking back at the wording, the first pick counts as one of the three queens

OpenStudy (kashmoneyjr):

So, you can throw out 1/49, which leaves you with 3/51*2/50

OpenStudy (naythyngo):

so it would be 3/51 times 2/51 which would give me 6/ 2550?

OpenStudy (kashmoneyjr):

2/50, since you lose a card by picking a queen

OpenStudy (kashmoneyjr):

but yeah 6/2550 sounds right

OpenStudy (naythyngo):

ok. thanks. do you think you can help me with another one. I haven't fully grasped the concept yet

OpenStudy (kashmoneyjr):

Sure!

OpenStudy (naythyngo):

Two number cubes are rolled for two separate events: All of the combinations of numbers on both cubes that give a sum less than 10. All combinations of numbers on both cubes that give a sum that is a multiple of 3.

OpenStudy (naythyngo):

i have to Complete the conditional-probability formula for event B given that event A occurs first by writing A and B in the blanks

OpenStudy (kashmoneyjr):

ok the formula is: P(B|A)=P(A and B)/P(A)

OpenStudy (kashmoneyjr):

and can be represented as an intersection, which is an upside down U

OpenStudy (kashmoneyjr):

the word "and" to be clear

OpenStudy (naythyngo):

i'm not sure how to fill in the equation though

OpenStudy (kashmoneyjr):

For this example?

OpenStudy (naythyngo):

yea with the number cubes

OpenStudy (kashmoneyjr):

number cubes? as in die?

OpenStudy (naythyngo):

I'm guessing but the question said number cube

OpenStudy (shelbywyatt):

isnt number cube same as die?

OpenStudy (naythyngo):

I'm pretty sure

OpenStudy (shelbywyatt):

so lets say it has number 1-6

OpenStudy (naythyngo):

ok.. how would i go about figuring out part a to this All of the combinations of numbers on both cubes that give a sum less than 10.

OpenStudy (shelbywyatt):

find every combinations that is a multiple of 3 and is less than 10

OpenStudy (naythyngo):

for the multiple of 3 would it just be 2/6

OpenStudy (shelbywyatt):

(1,2), (2,1), (4,2),(2,4),(3,3),(5,1),(1,5)(5,4)(4,5)(6,3)(3,6)

OpenStudy (shelbywyatt):

11/36

OpenStudy (naythyngo):

I'm a lot lost when it comes to answering the question

OpenStudy (shelbywyatt):

multiple choice?

OpenStudy (naythyngo):

no its a fill in.

OpenStudy (kashmoneyjr):

Could you provide the entire question?

OpenStudy (naythyngo):

Two number cubes are rolled for two separate events: All of the combinations of numbers on both cubes that give a sum less than 10. All combinations of numbers on both cubes that give a sum that is a multiple of 3. Complete the conditional-probability formula for event B given that event A occurs first by writing A and B in the blanks:

OpenStudy (kashmoneyjr):

So you basically have to consider the possibilities for each of these scenarios: Lets consider A: All of the combinations of numbers on both cubes that give a sum less than 10. We know that there are many less combinations that sum to >=10 than those that are less, so we can look at the ones that are greater or equal to. those combinations would be: (6,6), (6,6) (the other way), (6,5), (5,6), (6,4), (4,6), (5,5), (5,5) So those are 8/36 total combinations that are >=10. To find those that are <10, we simply subtract this value from 1, so 1-8/36=28/36

OpenStudy (kashmoneyjr):

We can use the formula to solve this, but it is much more effective to solve it logically

OpenStudy (naythyngo):

so how would I set it up for a conditional equation

OpenStudy (kashmoneyjr):

well you have to find P(A andB) and P(A)

OpenStudy (kashmoneyjr):

We found P(A), thats 28/36

OpenStudy (kashmoneyjr):

P(A and B) is the probability of getting a sum that is a multiple of 3 and is less than 10. Basically, the probability of getting a sum of 3, 6, or 9

OpenStudy (naythyngo):

9/36?

OpenStudy (kashmoneyjr):

So you have to find the total number of combinations within those a sum of 3 results from (1,2) or (2,1) A sum of 6 results from (1,5), (5,1), (4,2), (2,4), (3,3), and (3,3) 9 results from (3,6), (6,3), (4,5), (5,4) So I get 12/36 as P(A and B)

OpenStudy (kashmoneyjr):

So my calculations would then lead to (12/36)/(28/36), which is 12/28, or 3/7

OpenStudy (naythyngo):

idk how to fit the answers in to the blanks though

OpenStudy (kashmoneyjr):

the blanks are for P(A) and P(B)?

OpenStudy (naythyngo):

the blanks look weird

OpenStudy (naythyngo):

the blanks want me to fill in the equation

OpenStudy (kashmoneyjr):

well the equation is P(B|A)=P(A and B)/P(A) so, filled in, it would be: P(B|A)=(12/36)/(28/36)

OpenStudy (shelbywyatt):

hey hey but isnt 3,3 only one

OpenStudy (kashmoneyjr):

yes, but i would assume you count it twice since you can get it 2 ways, so it should be regarded as such when dealing with probability

OpenStudy (shelbywyatt):

oh ok

OpenStudy (shelbywyatt):

so we have to multiply 12/36 by 28/36?

OpenStudy (kashmoneyjr):

You would have to divide 12/36 by 28/36

OpenStudy (shelbywyatt):

6/18*14/18= 3/9*7/9

OpenStudy (shelbywyatt):

oh ok

OpenStudy (shelbywyatt):

3/9*9/7

OpenStudy (shelbywyatt):

3/1*1/7

OpenStudy (shelbywyatt):

3/7?

OpenStudy (kashmoneyjr):

Yes

OpenStudy (shelbywyatt):

wheres the owner of this forum